40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 1 12:12:35 2013, gautam, Update, CDS, New model for endtable PZTs C1ASX_GDS_TP.png
    Reply  Thu Aug 1 13:09:17 2013, gautam, Update, CDS, New model for endtable PZTs-procedure 
    Reply  Thu Aug 1 15:06:59 2013, jamie, Update, CDS, New model for endtable PZTs 
       Reply  Thu Aug 1 15:28:44 2013, gautam, Update, CDS, New model for endtable PZTs-problem solved 
    Reply  Thu Aug 1 20:58:56 2013, gautam, Update, CDS, New model for endtable PZTs-MEDM Screens setup MAIN_SCREEN.pdfMATRICES.pdfLOCKINS.pdfCONTROL_FILTERS.pdf
       Reply  Mon Aug 5 18:18:32 2013, gautam, Update, CDS, Choosing LO Amplitudes and Frequencies  GTR_Power_Spectrum.pdf
          Reply  Mon Aug 5 19:10:01 2013, Koji, Update, CDS, Choosing LO Amplitudes and Frequencies  
             Reply  Tue Aug 6 16:36:51 2013, gautam, Update, CDS, Choosing LO Amplitudes and Frequencies-revised powerSpec0806.pdf
                Reply  Wed Aug 7 23:40:49 2013, gautam, Update, CDS, X-End Green ASS - A first update time-series.pdf
                   Reply  Sat Aug 10 05:53:51 2013, gautam, Update, CDS, X-End Green ASS - Roundup New_ASX_MEDM_MAIN.pdf
                      Reply  Mon Aug 12 12:57:59 2013, Jenne, Update, CDS, X-End Green ASS - Roundup 
                         Reply  Mon Aug 12 13:30:33 2013, Jamie, Update, CDS, X-End Green ASS - Roundup 
Message ID: 8983     Entry time: Wed Aug 7 23:40:49 2013     In reply to: 8972     Reply to this: 8993
Author: gautam 
Type: Update 
Category: CDS 
Subject: X-End Green ASS - A first update 

 I have done some preliminary testing of the X-End Green ASS Servo. I will write a more detailed elog about this soon, but I thought I'd note down the important stuff here.


Yesterday, while we were venting, I aligned the X-arm to the green using the sliders on IFOalign, maximizing the transmission. Then I retook a power spectrum so as to determine the LO frequencies. Jenne pointed out that LO frequencies should not be integers (it usually suffices to append a .098725 or something to the frequency) so I made the necessary changes.

I did a first run of the servo yesterday, and more runs today. Notable points:

  1. I was able to lock to 00 from a 08 or 09 mode using the PZT sliders
  2. The green transmission having locked to 00 was ~0.2. I then ran the servo and got it up to ~0.4 and then 0.6 (see time series plot attached). The servo was able to recover this level of transmission after misaligning the steering mirrors using the PZT sliders.
  3. This was not the optimal transmission level as when Koji moved ETMX a little, the transmission improved.
  4. The actuators are degenerate. Most of the time, only two of the four servos are doing anything significant. This is probably because of the fact that the two steering mirrors are so close to each other, that moving one or the other produces virtually the same effect. I do however have some cool videos of mode-hopping :)
  5. The range of actuation of the PZTs is probably not enough to maximize the green transmission from an arbitrary state because of point 4 (i.e. we need to move one mirror in some direction a lot, and move the other a lot to compensate for the change, and the overall gain in input pointing/alignment is marginal). It may be that things will be slightly better at the Y-end. It would also be interesting to see if there is any improvement in the servo performance by dithering the cavity mirrors as opposed to the PZT mirrors.
  6. To this end, I tried modifying the c1asx model to incorporate an option to dither the cavity mirrors. The plan was to make a second set of LOs in the model that output to ITMX and ETMX suspensions. However, for some reason, when I recompiled the model and restarted it, c1iscex crashed. Parity has now been restored. Note that in order to accommodate the new LOs, I had to make changes to C1SUS, C1RFM and C1SCX as well. I have since removed all my additions, saved, built and installed these models, but have not restarted them (with the exception of C1SCX which restarted when I manually restarted c1iscex). 
  7. The plan tomorrow is to try incorporating cavity dither into the model again. This time, I'll try grabbing the LO-related signals from c1ass directly, as I am not clear why my approach did not work.

More details to follow.

time-series.pdf

ELOG V3.1.3-